Chelsea left the White House and traveled toward the capital at an average speed of 34 km/h. Jasmine left at the smae time and traveled in the opposite direction with an average speed of 65 km/h. Find the number of hours Jasmine needs to travel before theyh are 59.4 km apart.
Answer by mananth(16949)

chelsea - speed = 34 km/h
Jasmine - speed = 65 km/h
.
let jasmine travel a distance of x km when they are to be 59.4 km apart.
chelsea would travel 59.4 -x km
.
time taken by jasmine = x/34
time taken by chelsea = 59.4-x / 65
Since they start at same time the times are equal
.
x/34 = 59.4 -x / 65
65x=34(59.4-x)
65x=2019.6 - 34x
99x = 2019.6
x= 2019.6/99
x= 20.4 the distance traveled by Jasmine
Jasmines speed = 34km/h
Time taken = 20.4 /34
= 0.6 hours
Jasmine travels 0.6 hours